# Detection of Lensed Gravitational Waves from dark matter halos with deep learning

### v1

@@ Abstract
Context. Lensed gravitational waves (GWs) provide a new window into the study of dark matter substructures, yet the faint interference signatures they produce are buried in detector noise.
Aims. To address this challenge, we develop a deep learning framework based on a residual one dimensional convolutional neural network for lensed GW identification under multiband observations.
Methods. The model directly processes multiband waveforms from binary neutron star systems, covering the early inspiral observed by the DECi hertz Interferometer Gravitational wave Observatory (DECIGO) and the late inspiral observed by the Einstein Telescope (ET), which correspond approximately to the wave optics and geometrical optics regimes, respectively. It enables end to end classification of five classes: pure noise, unlensed GWs, and three representative lensed GWs corresponding to singular isothermal sphere (SIS), cored isothermal sphere (CIS), and Navarro–Frenk–White (NFW) profiles. A dataset of $10^{6}$ simulated samples was constructed with signal to noise ratios (SNR) from 5 to 100.
Results. The deep learning model with multiband observations achieves an accuracy of 97.0% and a macro averaged $F_{1}$ score of 0.97, significantly exceeding the single detector performance, where DECIGO and ET reach 72.8% and 62.3%, respectively. Even in the low SNR regime ($\mathrm{SNR}<20$) the model maintains an accuracy above 63%, while in the high SNR regime ($\mathrm{SNR}>80$) its accuracy approaches 99.8%.
Conclusions. These results demonstrate that multiband GW observations effectively enhance the detection of lensed GWs within complex noise environments, providing a robust and efficient pathway for the automated identification of lensed GWs in future multiband observations.
Key Words.:
lensed gravitational waves – dark matter – deep learning
1 Introduction
@@
Recent theoretical and numerical studies have shown that lensed GWs introduce distinctive amplitude modulations and time delays in the frequency domain, providing a unique probe of dark matter halos and primordial black holes (tambalo2023lensing; leung2025wave). Building on these findings, upcoming third generation detectors such as the Einstein Telescope (ET) and the space based DECi hertz Interferometer Gravitational wave Observatory (DECIGO) are expected to enable multiband observations that further enhance the detectability and parameter estimation precision of lensed GWs (abe2025multi; mukherjee2020multimessenger; li2023targeted). However, identifying and characterising lensed GWs remain challenging. Traditional matched filtering and Bayesian inference have played a central role in gravitational wave detection and parameter estimation, and continue to provide reliable results across various conditions. As observational sensitivity increases and the volume of detected events grows, these approaches may face practical limitations related to computational efficiency and waveform modeling (janquart2021fast; li2018gravitational). In particular, low SNR conditions and interference induced modulations can further complicate signal recovery. These developments highlight the importance of exploring complementary and adaptive analysis frameworks for next generation gravitational wave data.
@@
With the rapid development of deep learning and machine learning techniques, these methods have been progressively introduced into gravitational wave data analysis and lensed GWs identification (kim2021identification; goyal2021rapid; li2025identification; liu2025identifying). Compared with traditional matched filtering and Bayesian inference approaches, deep learning offers advantages in automatic feature extraction and efficient inference. Convolutional neural networks (CNNs) are capable of learning informative representations from time domain or time frequency domain data, which can help mitigate the dependence on predefined waveform templates. Moreover, their inference phase is computationally efficient, making them particularly suitable for real time identification and filtering of the massive number of GWs events expected in future observations (singh2018classifying; goyal2021rapid). Recent studies have demonstrated the potential of deep learning in identifying lensed GWs (wilde2022detecting), further confirming its applicability under various conditions. The advent of third generation detectors has also introduced new opportunities to overcome current limitations through multiband observations. Space based detectors such as the DECIGO exhibits higher sensitivity to low frequency wave optics interference effects, while ground based detectors such as the ET offer superior resolution for geometric optics magnification and phase delay phenomena at mid to high frequencies (abe2025multi). This complementarity across frequency bands implies that, by fusing data from multiple detectors, one can simultaneously exploit multiple scale physical features including diffraction and geometric magnification—to mitigate parameter degeneracies and enhance the accuracy of lensed GWs identification. A deep learning based, multiple detector framework for lensed GWs recognition is therefore expected to play a crucial role in future third generation and space based GWs detectors, providing new methodologies for probing dark matter structures and constraining cosmological parameters.
@@
The paper is structured as follows. Section 1 outlines the scientific motivation and background. Section 2 introduces the theoretical framework of lensed GWs and data generation. Section 3 describes the detector configuration and noise modeling. Section 4 presents the dataset construction and simulation setup. Section 5 details the deep learning architecture and training strategy. Section 6 reports the results and analyses under different detector configurations and SNR. Section 7 concludes with a summary and future outlook.
2 Theoretical Framework
@@
2.1 Fundamental Theory of Lensed GWs
@@
where $a$ is the cosmological scale factor and $h_{\mu\nu}$ is the dimensionless tensor perturbation representing the GWs disturbance. The Newtonian potential of the lens, $\Phi(\mathbf{x})$, satisfies $\Phi\ll 1$ to ensure the validity of the linear approximation. Under this condition, the GWs propagation equation can be approximated as a scalar wave equation with a potential term (abe2025multi):
@@
where $h(\tau,\mathbf{x})$ is the scalar form of the perturbation. Transforming into the frequency domain yields a Helmholtz-like equation (peters1974index):
@@
To quantify the lensing modulation, the complex amplification factor is defined (nakamura1999wave; takahashi2003wave):
@@
where $\tilde{h}_{\mathrm{UL}}$ denotes the unlensed waveform. Under the thin lens and Born approximations, $F$ can be expressed as a two dimensional integral over the lens plane (nakamura1999wave; takahashi2003wave):
@@
Hence, $w$ reflects the ratio between the Schwarzschild radius of the lens and the wavelength of the gravitational wave. A large $w$ corresponds to the geometrical optics regime ($w\gg 1$), while a small $w$ corresponds to the wave optics regime ($w\ll 1$) (abe2025multi). In this context, DECIGO mainly operates in the wave optics regime due to its lower frequency, whereas ET approaches the geometrical optics regime at higher frequencies.
@@
2.2 Lens Models
To investigate the lensed amplification effects of dark matter halos on GWs, three spherically symmetric mass density models are considered: the Singular Isothermal Sphere, the Cored Isothermal Spher, and the Navarro–Frenk–White profile. These models represent a gradual transition from dense to diffuse halo structures, spanning typical mass distributions from galactic to cluster scales. To focus on the dominant physical effects and maintain analytical tractability, all lenses are assumed to be perfectly spherical, neglecting secondary perturbations and substructure effects (poon2025galaxy). The analysis is performed under the thin lens and weak field approximations, allowing the amplification factor $F(f,y)$ to be derived in a unified formalism.
@@ 2.2 Lens Models
The frequency domain amplification factor is expressed analytically as a hypergeometric series (matsunaga2006finite):
@@
The CIS model extends the SIS profile by introducing a finite core radius $r_{c}$, representing halos with flattened central densities (kormann1994isothermal; flores1996cluster; treu2010strong). Its three dimensional and surface densities are
@@
The NFW profile, derived from cosmological $N$ body simulations, describes cold dark matter halos (navarro1996structure; navarro1997universal):
@@
In this work, the amplification factors $F(w,y)$ for the three representative dark matter halo lens models (SIS, CIS, and NFW)were computed using the GLoW_public library (GLoW_public).
3 Detector Configuration and Noise Modeling
This section outlines the detector configurations and noise models adopted in this study. It first introduces the representative detector systems and their sensitivity characteristics, followed by discussions of the antenna response functions (Sec. 3.2) and the modelling of detector noise (Sec. 3.3). A unified specification of these components ensures geometric and noise consistency between the DECIGO and the ET, establishing a consistent basis for waveform simulation and signal analysis of lensed GWs.
3.1 Detector Systems and Sensitivity Characteristics
This study considers two representative next generation GWs detectors: the space based DECIGO (yagi2011detector; kawamura2021current; tsuji2024significance) and the ground based ET (regimbau2012mock; punturo2010einstein; hild2009xylophone). These instruments are complementary in sensitivity and operational frequency range, covering approximately $0.1$-$10\penalty 10000\ \mathrm{Hz}$ for DECIGO and $10$-$1000\penalty 10000\ \mathrm{Hz}$ for ET. The inspiral phase of BNS merger spans frequencies from $0.1$ to $10^{3}\penalty 10000\ \mathrm{Hz}$, tracing the system’s evolution from early orbital contraction to the pre-merger stage. DECIGO is thus sensitive to the low frequency early inspiral, whereas ET captures the mid to high frequency band preceding coalescence. The noise power spectral density (PSD) for DECIGO follows the official design specification (yagi2011detector), and that for ET is based on the ET sensitivity curve (hild2011sensitivity).
Fig. 1 compares the PSD of DECIGO and ET with the amplitude spectral density (ASD) of a typical BNS signal. DECIGO achieves higher sensitivity in the $0.1$-$10\penalty 10000\ \mathrm{Hz}$ band, enabling the detection of interference and diffraction features produced by low mass lenses dominated by wave optics effects. In contrast, ET exhibits superior response in the $20$-$200\penalty 10000\ \mathrm{Hz}$ range, where geometric optics amplification and phase shifts prevail. The complementarity of their frequency coverage allows the same merger event to be jointly observed during different inspiral stages, yielding a more complete waveform and enhancing the accuracy of lens type classification.
@@ 3.1 Detector Systems and Sensitivity Characteristics
3.2 Detector Response Functions
@@
DECIGO is a space based interferometer with a triangular three arm configuration whose antenna response varies periodically with time. Let $(\bar{\theta}_{S},\bar{\phi}_{S})$ denote the source position in the heliocentric coordinate system and $(\hat{\theta}_{L},\hat{\phi}_{L})$ the orientation of the orbital angular momentum. The observed strain in the detector frame is (yagi2011detector)
@@
where $T=1\penalty 10000\ \mathrm{yr}$ is the orbital period and $c_{0}=0$ is the initial orbital phase.
The Einstein Telescope consists of three co-located interferometers arranged in an equilateral triangle, each having equal arm lengths and an opening angle of $\gamma=\pi/3$. Let $(\theta,\phi)$ denote the source sky location in the detector frame and $\psi$ the polarization angle. For a single interferometer with arm orientation offset $\Delta$, the antenna pattern functions for the plus and cross polarizations are given by (hild2011sensitivity)
@@
3.3 Noise Modeling and Detector Noise Generation
Based on the one sided noise PSD, we could obtain the time domain noise signal from the one sided PSD; in this paper, we utilized the Python function pycbc.noise.gaussian.noise_from_psd (alex_nitz_2024_10473621), which takes a PSD as input and returns colored Gaussian noise, to simulate the time domain noise received by both the DECIGO and ET detectors. The corresponding PSDs of ET and DECIGO were adopted according to their official sensitivity curves (yagi2011detector; hild2011sensitivity), ensuring that the generated noise realizations accurately reflect the instrumental characteristics across their respective frequency bands.
4 Dataset Construction and Simulation Setup
@@
Figure 2: Frequency and time domain waveforms of BNS ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100\penalty 10000\ \mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by DECIGO. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
@@
Figure 3: Frequency and time domain waveforms of ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100\penalty 10000\ \mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by ET. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
@@
Primary mass $m_{1}$ $\mathcal{U}(1,3)\penalty 10000\ M_{\odot}$
Secondary mass $m_{2}$ $\mathcal{U}(1,3)\penalty 10000\ M_{\odot}$
Luminosity distance $D_{L}$ $\mathcal{U}(0,10^{3})\penalty 10000\ \mathrm{Mpc}$
@@
For both the space based DECIGO and the ground based ET detectors, five classes of samples were generated: pure noise (PN), unlensed (UL), and three lensed cases (SIS, CIS, NFW). The physical parameters of the BNS, including component masses and luminosity distances, were randomly and uniformly sampled according to Table 1. To evaluate the impact of SNR on classification performance, the data were stratified by the SNR of the UL signal in noise, divided into five intervals: 55-2020, 2020-$40$, 4040-$60$, 6060-$80$, and 8080-$100$. Each SNR interval contained an independent dataset comprising five signal classes (PN, UL, SIS, CIS, NFW) with $4\times 10^{4}$ samples per class, yielding $2\times 10^{5}$ samples per set. All signals shared the same parameter sampling scheme but differed in noise realizations, lens classes, and modulation parameters. For each detector, datasets were generated following identical SNR intervals and sample counts, ensuring direct comparability between DECIGO and ET under matched physical conditions. The definitions of all sample classes are listed in Table 2.
After generating the SNR stratified datasets, the data were further organized for different experimental configurations. For each detector (ET or DECIGO), all SNR intervals were merged into a unified full range dataset spanning 55-$100$, comprising roughly $10^{6}$ samples, to train single detector models and evaluate overall classification performance. To enable a fair comparison with models trained on individual detectors, a model was trained on the combined dataset from the multiple detector configuration (ET+DECIGO), within the same SNR range. This approach allows us to evaluate the potential performance gains introduced by multiband observations. Additionally, to investigate the effect of SNR, we trained and evaluated separate models across five distinct SNR intervals.
DECIGO operates within the $0.1$-$10\penalty 10000\ \mathrm{Hz}$ band, with a sampling rate of $1\penalty 10000\ \mathrm{Hz}$ and signal duration of $3600\penalty 10000\ \mathrm{s}$, while ET covers $20$-$200\penalty 10000\ \mathrm{Hz}$ with a sampling rate of $128\penalty 10000\ \mathrm{Hz}$ and duration of $20\penalty 10000\ \mathrm{s}$. The waveform templates were generated based on the post-Newtonian model described in Ref. wu2024comparison; blanchet2024post; blanchet2019analyzing; cho2022gravitational, adopting a 2.52.5PN approximation including spin–orbit (SO) coupling while neglecting spin–spin (SS) effects. The BNS component masses and luminosity distance were sampled according to Table 2, while the remaining source parameters were fixed at inclination $\iota=1.0$, initial phase $\phi_{0}=0$, and dimensionless spins $\chi_{1}=\chi_{2}=0.1$. The amplification factor $F(w,y)$ was computed numerically using the GLoW_public (GLoW_public), with fixed lens parameters $M_{L}=2\times 10^{3}\,M_{\odot}$, $y=0.3$, $x_{c}=0.3$, and $x_{s}=0.3$, where $M_{L}$ is the lens mass, $y$ is the dimensionless impact parameter, and $x_{c}$ and $x_{s}$ are the dimensionless core radii for the CIS and NFW models, respectively.
For the adopted parameters, the characteristic time delay between multiple images can be estimated as(abe2025multi)
$\Delta t_{\mathrm{d}}\simeq 0.03\left(\frac{M_{Lz}}{3\times 10^{3}M_{\odot}}\right)\left(\frac{y}{0.3}\right)\penalty 10000\ \mathrm{s}.$ (39)
@@
Class frequency domain form time domain form
@@
SIS $F_{\rm SIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm SIS}(t)+n(t)$
CIS $F_{\rm CIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm CIS}(t)+n(t)$
NFW $F_{\rm NFW}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm NFW}(t)+n(t)$
@@
5 Deep Learning Architecture and Training Strategy

### v2

# Detection of Multiband Lensed Gravitational Waves from Dark Matter Halos with Deep Learning

@@ Abstract
Lensed gravitational waves (GWs) propagate through perturbed gravitational potentials and produce interference patterns in amplitude and phase. These patterns contain information about dark matter substructures, yet their signatures are often buried in detector noise. To address this challenge, we develop a deep learning framework for identifying lensed signals under multiband observations. The model processes binary neutron star waveforms, with the early inspiral observed by the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) and the late inspiral captured by the Einstein Telescope (ET). These two bands approximately correspond to the wave optics and geometric optics regimes. We consider five classes of data: pure noise, unlensed signals, and three lensed signals generated by the Singular Isothermal Sphere, Cored Isothermal Sphere, and Navarro–Frenk–White profiles. Using a dataset of $10^{6}$ samples with signal to noise ratio values from 5 to 100, the model achieves an accuracy of $97.0\%$, notably exceeding the performance of single detector configurations such as DECIGO or ET. The model maintains an accuracy above $63\%$ when the SNR is below $20$, and the accuracy increases to $99.8\%$ once the SNR exceeds $80$. These results demonstrate that multiband observations improve the detectability of lensing effects and provide an efficient and robust approach for automated identification of lensed GWs.
preprint: APS/123-QED
I Introduction
@@
Recent theoretical and numerical studies have shown that lensed GWs introduce distinctive amplitude modulations and time delays in the frequency domain, providing a unique probe of dark matter halos and primordial black holes [tambalo2023lensing, leung2025wave]. Building on these findings, upcoming third generation detectors such as the Einstein Telescope (ET) and the space based DECi hertz Interferometer Gravitational wave Observatory (DECIGO) are expected to enable multiband observations that further enhance the detectability and parameter estimation precision of lensed GWs [abe2025multi, mukherjee2020multimessenger, li2023targeted]. However, identifying and characterising lensed GWs remain challenging. Traditional matched filtering and Bayesian inference have played a central role in gravitational wave detection and parameter estimation, and continue to provide reliable results across various conditions. As observational sensitivity increases and the volume of detected events grows, these approaches may face practical limitations related to computational efficiency and waveform modeling [janquart2021fast, li2018gravitational]. In particular, low SNR conditions and interference induced modulations can further complicate signal recovery. These developments highlight the importance of exploring complementary and adaptive analysis frameworks for next generation gravitational wave data.
@@
With the rapid development of deep learning and machine learning techniques, these methods have been progressively introduced into gravitational wave data analysis and lensed GWs identification [kim2021identification, goyal2021rapid, li2025identification, liu2025identifying, chan2025detectability, chan2025identification]. Compared with traditional matched filtering and Bayesian inference approaches, deep learning offers advantages in automatic feature extraction and efficient inference. Convolutional neural networks (CNNs) are capable of learning informative representations from time domain or time frequency domain data, which can help mitigate the dependence on predefined waveform templates. Moreover, their inference phase is computationally efficient, making them particularly suitable for real time identification and filtering of the massive number of GWs events expected in future observations [singh2018classifying, goyal2021rapid]. Recent studies have demonstrated the potential of deep learning in identifying lensed GWs [wilde2022detecting], further confirming its applicability under various conditions. The advent of third generation detectors has also introduced new opportunities to overcome current limitations through multiband observations. Space based detectors such as the DECIGO exhibits higher sensitivity to low frequency wave optics interference effects, while ground based detectors such as the ET offer superior resolution for geometric optics magnification and phase delay phenomena at mid to high frequencies [abe2025multi]. This complementarity across frequency bands implies that, by fusing data from multiple detectors, one can simultaneously exploit multiple scale physical features including diffraction and geometric magnification—to mitigate parameter degeneracies and enhance the accuracy of lensed GWs identification. A deep learning based, multiple detector framework for lensed GWs recognition is therefore expected to play a crucial role in future third generation and space based GWs detectors, providing new methodologies for probing dark matter structures and constraining cosmological parameters.
@@
The paper is structured as follows. Section I outlines the scientific motivation and background. Section II introduces the theoretical framework of lensed GWs and data generation. Section III describes the detector configuration and noise modeling. Section IV presents the dataset construction and simulation setup. Section V details the deep learning architecture and training strategy. Section VI reports the results and analyses under different detector configurations and SNR. Section VII concludes with a summary and future outlook.
II Theoretical Framework
@@
II.1 Fundamental Theory of Lensed GWs
@@
where $a$ is the cosmological scale factor and $h_{\mu\nu}$ is the dimensionless tensor perturbation representing the GWs disturbance. The Newtonian potential of the lens, $\Phi(\mathbf{x})$, satisfies $\Phi\ll 1$ to ensure the validity of the linear approximation. Under this condition, the GWs propagation equation can be approximated as a scalar wave equation with a potential term [abe2025multi]:
@@
where $h(\tau,\mathbf{x})$ is the scalar form of the perturbation. Transforming into the frequency domain yields a Helmholtz-like equation [peters1974index]:
@@
To quantify the lensing modulation, the complex amplification factor is defined [nakamura1999wave, takahashi2003wave]:
@@
where $\tilde{h}_{\mathrm{UL}}$ denotes the unlensed waveform. Under the thin lens and Born approximations, $F$ can be expressed as a two dimensional integral over the lens plane [nakamura1999wave, takahashi2003wave]:
@@
Hence, $w$ reflects the ratio between the Schwarzschild radius of the lens and the wavelength of the gravitational wave. A large $w$ corresponds to the geometrical optics regime ($w\gg 1$), while a small $w$ corresponds to the wave optics regime ($w\ll 1$[abe2025multi]. In this context, DECIGO mainly operates in the wave optics regime due to its lower frequency, whereas ET approaches the geometrical optics regime at higher frequencies.
@@
II.2 Lens Models
To investigate the lensed amplification effects of dark matter halos on GWs, three spherically symmetric mass density models are considered: the Singular Isothermal Sphere, the Cored Isothermal Spher, and the Navarro–Frenk–White profile. These models represent a gradual transition from dense to diffuse halo structures, spanning typical mass distributions from galactic to cluster scales. To focus on the dominant physical effects and maintain analytical tractability, all lenses are assumed to be perfectly spherical, neglecting secondary perturbations and substructure effects [poon2025galaxy]. The analysis is performed under the thin-lens and weak field approximations, allowing the amplification factor $F(f,y)$ to be derived in a unified formalism.
@@ II.2 Lens Models
The frequency domain amplification factor is expressed analytically as a hypergeometric series [matsunaga2006finite]:
@@
The CIS model extends the SIS profile by introducing a finite core radius $r_{c}$, representing halos with flattened central densities [kormann1994isothermal, flores1996cluster, treu2010strong]. Its three dimensional and surface densities are
@@
The NFW profile, derived from cosmological $N$ body simulations, describes cold dark matter halos [navarro1996structure, navarro1997universal]:
@@
In this work, the amplification factors $F(w,y)$ for the three representative dark matter halo lens models (SIS, CIS, and NFW)were computed using the GLoW_public library [GLoW_public].
III Detector Configuration and Noise Modeling
This section outlines the detector configurations and noise models adopted in this study. It first introduces the representative detector systems and their sensitivity characteristics, followed by discussions of the antenna response functions (Sec. III.2) and the modelling of detector noise (Sec. III.3). A unified specification of these components ensures geometric and noise consistency between the DECIGO and the ET, establishing a consistent basis for waveform simulation and signal analysis of lensed GWs.
III.1 Detector Systems and Sensitivity Characteristics
[…]
@@ III.1 Detector Systems and Sensitivity Characteristics
III.2 Detector Response Functions
@@
DECIGO is a space based interferometer with a triangular three arm configuration whose antenna response varies periodically with time. Let $(\bar{\theta}_{S},\bar{\phi}_{S})$ denote the source position in the heliocentric coordinate system and $(\hat{\theta}_{L},\hat{\phi}_{L})$ the orientation of the orbital angular momentum. The observed strain in the detector frame is [yagi2011detector]
@@
where $T=1~\mathrm{yr}$ is the orbital period and $c_{0}=0$ is the initial orbital phase.
The Einstein Telescope consists of three co-located interferometers arranged in an equilateral triangle, each having equal arm lengths and an opening angle of $\gamma=\pi/3$. Let $(\theta,\phi)$ denote the source sky location in the detector frame and $\psi$ the polarization angle. For a single interferometer with arm orientation offset $\Delta$, the antenna pattern functions for the plus and cross polarizations are given by [hild2011sensitivity]
@@
III.3 Noise Modeling and Detector Noise Generation
Based on the one sided noise PSD, we could obtain the time domain noise signal from the one sided PSD; in this paper, we utilized the Python function pycbc.noise.gaussian.noise_from_psd [alex_nitz_2024_10473621], which takes a PSD as input and returns colored Gaussian noise, to simulate the time domain noise received by both the DECIGO and ET detectors. The corresponding PSDs of ET and DECIGO were adopted according to their official sensitivity curves [yagi2011detector, hild2011sensitivity], ensuring that the generated noise realizations accurately reflect the instrumental characteristics across their respective frequency bands.
IV Dataset Construction and Simulation Setup
@@
Figure 2: Frequency and time domain waveforms of BNS ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100~\mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by DECIGO. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
@@
Figure 3: Frequency and time domain waveforms of ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100~\mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by ET. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
@@
Primary mass $m_{1}$ $\mathcal{U}(1,3)~M_{\odot}$
Secondary mass $m_{2}$ $\mathcal{U}(1,3)~M_{\odot}$
Luminosity distance $D_{L}$ $\mathcal{U}(0,10^{3})~\mathrm{Mpc}$
@@
For both the space based DECIGO and the ground based ET detectors, five classes of samples were generated: pure noise (PN), unlensed (UL), and three lensed cases (SIS, CIS, NFW). The physical parameters of the BNS, including component masses and luminosity distances, were randomly and uniformly sampled according to Table 1. To evaluate the impact of SNR on classification performance, the data were stratified by the SNR of the UL signal in noise, divided into five intervals: 552020, 20204040, 40406060, 60608080, and 8080100100. Each SNR interval contained an independent dataset comprising five signal classes (PN, UL, SIS, CIS, NFW) with $4\times 10^{4}$ samples per class, yielding $2\times 10^{5}$ samples per set. All signals shared the same parameter sampling scheme but differed in noise realizations, lens classes, and modulation parameters. For each detector, datasets were generated following identical SNR intervals and sample counts, ensuring direct comparability between DECIGO and ET under matched physical conditions. The definitions of all sample classes are listed in Table 2.
After generating the SNR stratified datasets, the data were further organized for different experimental configurations. For each detector (ET or DECIGO), all SNR intervals were merged into a unified full range dataset spanning 55100100, comprising roughly $10^{6}$ samples, to train single detector models and evaluate overall classification performance. To enable a fair comparison with models trained on individual detectors, a model was trained on the combined dataset from the multiple detector configuration (ET+DECIGO), within the same SNR range. This approach allows us to evaluate the potential performance gains introduced by multiband observations. Additionally, to investigate the effect of SNR, we trained and evaluated separate models across five distinct SNR intervals.
[…]
For the adopted parameters, the characteristic time delay between multiple images can be estimated as[abe2025multi]
$\Delta t_{\mathrm{d}}\simeq 0.03\left(\frac{M_{Lz}}{3\times 10^{3}M_{\odot}}\right)\left(\frac{y}{0.3}\right)~\mathrm{s}.$ (39)
@@
Class frequency domain construction time domain form
@@
SIS lens $F_{\rm SIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm SIS}(t)+n(t)$
CIS lens $F_{\rm CIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm CIS}(t)+n(t)$
NFW lens $F_{\rm NFW}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm NFW}(t)+n(t)$
@@
V Deep Learning Architecture and Training Strategy